# Kilohertz Gravitational Waves from Binary Neutron Star Mergers: Full Spectrum Analyses and High-density Constraints on Neutron Star Matter

### v1

@@ Abstract
We demonstrate Bayesian analyses of the complete gravitational-wave spectrum of binary neutron star mergers events with the next-generation detector Einstein Telescope. Our mock analyses are performed for 20 different signals using the TEOBResumSPA_NRPMw waveform that models gravitational-waves from the inspiral to the postmerger phase. They are employed to validate a pipeline for neutron star’s extreme matter constraints with prospective detections and under minimal hypotheses on the equation of state. The proposed analysis stack delivers inferences for the mass-radius curve, the mass dependence of the quadrupolar tidal polarizability parameter, the neutron star’s maximum density, the maximum mass and the relative radius, and the pressure-density relation itself. We show that a single event at a signal-to-noise ratio close to the minimum threshold for postmerger detection is sufficient to tightly constrain all the above relations as well as quantities like the maximum mass (maximum density) to precision of ${\sim}6$% (${\sim}10$%) at 90% credibility level. We also revisit inferences of prompt black hole formation with full spectrum signals and find that the latter can be robustly identified, even when the postmerger is not detectable due to a low signal-to-noise ratio. New results on the impact of the initial signal frequency and of the detector configuration (triangular vs. two-L) on the source’s parameters estimation are also reported. An improvement of approximately one order of magnitude in the precision of the chirp mass and mass ratio can be achieved by lowering the initial frequency from 20 Hz to 2 Hz. The two-L configuration shows instead significant improvements on the inference of the source declination, due to geographical separation of the two detectors.
@@ I Introduction
Binary neutron stars are the main sources for next-generation (XG) ground-based gravitational wave (GW) detectors, such as Einstein Telescope (ET) Hild et al. (2011); Hild (2012); Abac et al. (2025); Maggiore et al. (2020); Punturo et al. (2010a) and Cosmic Explorer (CEReitze et al. (2019); Evans et al. (2021). They are expected to be observed with a rate of $7\times 10^{4}$ mergers per year Belgacem et al. (2019); Kalogera et al. (2021); Iacovelli et al. (2022). ET is designed with a xylophone configuration, combining a cryogenic interferometer for the low-frequency band with a high-power interferometer targeting frequencies above ${\sim}100~{}{\rm Hz}$. This design will not only significantly enhance current GW detector sensitivity, but also enable the observations of postmerger (PM) signals at kiloHertz frequencies. GWs observations of the full spectrum (inspiral-merger-postmerger) of a BNS merger can convey unique information on the nuclear matter that constitutes this compact object e.g. (Abbott et al., 2018a; Bauswein et al., 2017; Breschi et al., 2019; Radice et al., 2020; Breschi et al., 2022a; Wijngaarden et al., 2022), thus improving current GW constraints after GW170817 Abbott et al. (2017a, b, 2018b). While the inspiral-merger (IM) carries the footprint of the matter properties at the progenitor densities, i.e. $\rho\lesssim{2}\,\rho_{\rm sat}$ (where $\rho_{\rm sat}\simeq 2.7{\times}10^{14}~{}{\rm g}~{}{\rm cm}^{-3}$ is the nuclear saturation density), the kiloHertz PM transient encodes information about the high-density (${\sim}3-6\rho_{\rm sat}$) regime of the equation of state (EoSBreschi et al. (2022a), enabling the investigation of unrevealed properties of the nuclear matter at high-densities, unreachable with modern experiments Radice et al. (2017); Bauswein et al. (2019); Breschi et al. (2022a); Most and Raithel (2021); Prakash et al. (2021).
@@
Inference of source properties and constraints on the neutron star (NS) EoS requires the availability of complete waveform templates for matched-filtering analyses. These templates are crucially designed using empirical relations computed from numerical relativity (NR) simulations and connecting morphological features of the PM signal to NS properties, e.g. Shibata and Uryu (2002); Stergioulas et al. (2011); Bauswein and Janka (2012); Bauswein et al. (2012); Hotokezaka et al. (2013); Takami et al. (2014); Bauswein and Stergioulas (2015); Bernuzzi et al. (2015); Breschi et al. (2022a, 2024a). In particular, Breschi et al. (2019) proposed the first time-domain full spectrum model that was constructed by combining an effective-one-body (EOB) IM waveform with a NR-informed analytical PM model (see also Breschi et al. (2024a, 2022b)). Similar approaches have been used in Wijngaarden et al. (2022); Puecher et al. (2023a), where phenomenological frequency-domain IM waveforms are stitched to PM models based on wavelets or Lorenzian templates. In all cases the phenomenological PM parameters are connected to the source parameter using NR empirical relations; see also (Bauswein et al., 2016; Bose et al., 2018; Soultanis et al., 2022; Vretinaris et al., 2025) for other examples of analytical PM models and (Easter et al., 2020; Whittaker et al., 2022) for data-driven PM models, which have not been incorporated in full spectrum models yet.
@@
These modeling efforts, combined with the exceptional sensitivity of XG detectors, provide a crucial foundation for accurate EoS inference. Notably, even a single detection of BNS with ET may suffice to place stringent constraints on the NS mass-radius relation Breschi et al. (2022a). That approach enabled tight constraints of the mass-radius diagram and demonstrated that the maximum NS mass can be determined with an accuracy better than 12%. Additional works Breschi et al. (2023); Rezzolla and Takami (2016); Chatziioannou et al. (2017); Lehoucq et al. (2025); Vretinaris et al. (2025) have similarly combined information from the IM section to PM features, highlighting the complementarity of these two regimes in probing supranuclear matter. Numerous proof of principle studies focus exclusively on the PM GW emission from BNSs to investigate the high-density nuclear matter through NR simulations. These works exploit the connection between peak frequencies and EoS-dependent properties, such as total mass and compactness Bauswein et al. (2014); Most and Raithel (2021); Easter et al. (2019); Takami et al. (2015); Bauswein et al. (2016); Ecker et al. (2025); Bamber et al. (2025); Mitra et al. (2025). Furthermore, recent works explored the potential for EoS inference using advanced and XG detectors Walker et al. (2024); Gupta et al. (2022); Iacovelli et al. (2023); Bandopadhyay et al. (2024); Kiendrebeogo et al. (2023) and assess the impact of different waveform models on nuclear EoS inference Yelikar et al. (2024).
@@
In this paper, we perform an extensive investigation of the capabilities of ET to infer BNS properties. We firstly focus on the low-frequency regime, assessing the impact on source parameter estimation of the initial frequency and of detector configurations (two-L and triangular). We then concentrate in the detections of the complete GW spectrum and study the prospective accuracy of EoS inference. We develop and employ a full spectrum model which combines the fast frequency-domain EOB TEOBResumSPA Akcay et al. (2019); Gamba et al. (2021a, 2023) with the frequency-domain wavelet-based NRPMw model Breschi et al. (2024a, 2022b) (Paper I and Paper II hereafter). We then perform Bayesian mock analyses of 20 fiducial BNSs, covering the frequency range from $f_{0}=20$ Hz up to the kiloHertz regime. These analyses serve not only to demonstrate the feasibility of combining information across the full GW spectrum, from inspiral to merger and postmerger, but also to quantify the achievable accuracy in the inference of source parameters. Beyond parameter estimation, our primary objective is to assess the improvements in constraints on the supranuclear properties governing NS matter. By incorporating both low- and high-frequency components of the signal, we explore how analyses of ET signals enhance our ability to probe the dense-matter regime, including tighter bounds on the mass-radius relation, the maximum mass supported by a stable non-rotating NS and the PC inference.
@@
All quantities are expressed in SI units, with masses in Solar masses ${\rm M_{\odot}}$ and distances in Mpc. The symbol ‘$\log$’ denotes the natural logarithm. The waveform strain is decomposed in multipoles $(\ell,m)$ as
@@
where $D_{L}$ is the luminosity distance, ${}_{-2}Y_{\ell m}$ are the $s=-2$ spin-weighted spherical harmonics and $\iota,\varphi$ are respectively the polar and azimuthal angles that define the orientation of the binary with respect to the observer. Each waveform mode $h_{\ell m}(t)$ is decomposed in amplitude $A_{\ell m}(t)$ and phase $\phi_{\ell m}(t)$, as
@@
The waveform depends on intrinsic and extrinsic source parameters. The intrinsic parameters of a BNS system are masses, spins and tidal polarizability parameters, e.g. $\boldsymbol{\theta}_{\rm bin}=\{M,q,\chi_{1},\chi_{2},\Lambda_{1},\Lambda_{2}\}$. The total binary mass is indicated with $M=m_{1}+m_{2}$, the mass ratio is $q=m_{1}/m_{2}\geq 1$, and the symmetric mass ratio is $\nu=m_{1}m_{2}/M^{2}$. The dimensionless spin vectors are denoted with ${\boldsymbol{\chi}}_{i}$ for $i=1,2$ and the spin component aligned with the orbital angular momentum L are labeled as $\chi_{i}={\boldsymbol{\chi}}_{i}\cdot\textbf{L}/|\textbf{L}|$. The effective spin parameter ${\chi_{\rm eff}}$ is a mass-weighted, aligned spin combination, i.e.
@@
The quadrupolar tidal polarizability parameters are defined as $\Lambda_{i}=({2}/{3})\,k_{2,i}\,C_{i}^{-5}$ for $i=1,2$, where $k_{2,i}$ and $C_{i}$ are the second Love number and the compactness of the $i$-th object, respectively. The reduced tidal parameter $\tilde{\Lambda}$ is
@@
The extrinsic parameters of the source ${\boldsymbol{\theta}_{\rm ext}}=\{D_{L},\iota,\alpha,\delta,\psi,t_{\rm mrg},\phi_{\rm mrg}\}$, i.e. inclination angle $\iota$, right ascension angle $\alpha$, declination angle $\delta$, polarization angle $\psi$, time of coalescence $t_{\rm mrg}$, and phase at the merger $\phi_{\rm mrg}$, allow us to identify the location and orientation of the source. Moreover, we include the recalibration parameters $\boldsymbol{\delta}_{\rm fit}$ that account for deviations from the predictions of the EoS-insensitive relations consistently with the related theoretical uncertainties. NRPMw is parameterized by two additional sets of degrees of freedom (see Paper I for a detailed discussion). The PM parameters ${\boldsymbol{\theta}_{\rm PM}^{\rm free}}=\{\phi_{\rm PM},t_{c},\alpha_{\rm drift}\}$ correspond to: the PM phase-shift $\phi_{\rm PM}$ that identifies the phase discontinuity after merger; the time of collapse $t_{c}$ that characterizes the collapse of the remnant into BH; and the frequency drift $\alpha_{\rm drift}$ that accounts for linear evolution of the dominant $f_{2}$ component.
@@ II.1 Waveform model
Figure 1: Attachment between TEOBResumSPA and NRPMw for a GW template of non-spinning binary with $m_{1}=1.5~{}{\rm M_{\odot}}$, $m_{2}=1.4~{}{\rm M_{\odot}}$, $\Lambda_{1}=400$, $\Lambda_{2}=600$ and $D_{L}=40~{}{\rm Mpc}$. Since the model is defined in the frequency domain, we obtain $h_{22}(t)$ by performing an inverse FFT of $h_{22}(f)$.
@@
We attach the EOB model for the IM to the PM NRPMw template, focusing on the dominant $(2,2)$ mode. The information required to perform the attachment of NRPMw with TEOBResumSPA is the time and the phase at merger, respectively $t_{\rm mrg}$ and $\phi_{\rm mrg}$. These data can be extracted from the EOB IM waveform. Then, resorting linear properties of the Fourier transform, the inspiral-merger-postmerger (IMPM) waveform can be computed as the sum of the pre-merger template $h_{\rm EOB}$ and the PM one $h_{\rm PM}$ in the frequency-domain,
@@
We also validate the model against six NR simulations that are not included in the calibration set of NRPMw, finding that the model is $89-95\%$ faithful to NR when considering the full spectrum of frequencies and the ET-D sensitivity curve Hild et al. (2011); Hild (2012). More details on the IMPM model and its validation can be found in Appendix A.
@@ II.2 Inference
Figure 2: Amplitude spectral density (ASD) of ET in configuration D Hild et al. (2011); Hild (2012) (in black) and comparison with ASD of LIGO L1 detector, relative to GW170817 event Abbott et al. (2019a) (in grey), to show the improvement in the detection of GW signals with XG detectors. The solid colored lines represent five exemplary BNS waveform spectra with TEOBResumSPA_NRPMw template for $\mathcal{M}=1.24~{}{\rm M_{\odot}}$, $q=1$, varying $\tilde{\Lambda}$ and $D_{L}$ as indicated in the legend.
We perform Bayesian parameter estimation (PE) experiments by injecting and analyzing mock zero-noise BNS signals with XG detectors. We employ the power spectral density (PSD) curve of the ET detector, as reported in Fig. 2, in configuration D Hild et al. (2011); Hild (2012), i.e. an equilateral triangular shape with $10~{}{\rm km}$ long arms. We simulate both IM-only and IMPM waveforms and recover them using the same model employed for the injection, TEOBResumSPA and TEOBResumSPA_NRPMw respectively.
@@
Bayesian analysis is based on the evaluation of the likelihood function (See Appendix B.1 for additional details), which is the most computationally expensive part. To speed up the computations, we make use of the relative binning technique Dai et al. (2018); Zackay et al. (2018) in the IM part. Relative binning consists in the precomputation of frequency-binned overlap of a fiducial waveform with the data, since the ratio of neighboring waveforms in frequency domain is smooth in the inspiral part. In particular, for the IM-only PEs we employ the relative binning method for the whole frequency range of the integration. For the IMPM cases, we utilize relative binning only up to a cutoff frequency $f_{\rm cut}=2048~{}{\rm Hz}$ and we perform a normal Bayesian PE without speed up after the cutoff frequency. The two likelihoods are then combined by summing their respective contributions:
@@
where $\log p({\text{{d}}}_{\rm IM}|\boldsymbol{\theta},H_{\rm IM}^{\rm RB})$ is the likelihood evaluated in the IM portion of the spectrum with relative binning, while $\log p({\text{{d}}}_{\rm PM}|\boldsymbol{\theta},H_{\rm PM})$ is evaluated in the high-frequency section, i.e. with NRPMw. For further explanation of the relative binning implementation and a discussion of the approximations used, we refer to Appendix B.2.
II.3 EoS constraints from quasi-universal relations
Given an estimation of the intrinsic BNS parameters, we show how XG detectors can enable precise constraints on the EoS of NSs. Inference of IM delivers posteriors of the masses of the compact objects and of the tidal polarizability parameters. The information contained in the IMPM signals, instead, allows to go beyond $m_{i}$ and $\Lambda_{i}$ and probe high-density NS properties. Breschi et al. (2022a) combined the Fisher matrix approach in the IM part with the Bayesian analysis of the PM in order to obtain a tighter constraint on the mass-radius relation. Here we perform a full spectrum Bayesian PE, aiming to obtain more reliable inference, particularly in the high-density regime. The resulting pipeline constrains different quantities of interest in EoS modeling and is described in what follows.
The first step is to map the posteriors for the total mass and the peak PM frequency $f_{2}$ into posteriors of maximum density ${\rho_{\rm max}^{\rm TOV}}$ and radius at the maximum mass ${R_{\rm max}^{\rm TOV}}$, using the NR phenomenological relations Breschi et al. (2022a, 2024b):

### v2

@@ Abstract
We demonstrate Bayesian analyses of the complete gravitational-wave spectrum of binary neutron star mergers events with the next-generation detector Einstein Telescope. Our mock analyses are performed for 20 different signals using the TEOBResumSPA_NRPMw waveform that models gravitational waves from the inspiral to the postmerger phase. They are employed to validate a pipeline for neutron star’s extreme matter constraints with prospective detections and under minimal hypotheses on the equation of state. The proposed analysis stack delivers inferences for the mass-radius curve, the mass dependence of the quadrupolar tidal polarizability parameter, the neutron star’s maximum density, the maximum mass and the relative radius, and the pressure-density relation itself. We show that a single event at a signal-to-noise ratio close to the minimum threshold for postmerger detection is sufficient to tightly constrain all the above relations as well as quantities like the maximum mass (maximum density) to precision of ${\sim}6$% (${\sim}10$%) at 90% credibility level. We also revisit inferences of prompt black hole formation with full spectrum signals and find that the latter can be robustly identified, even when the postmerger is not detectable due to a low signal-to-noise ratio. New results on the impact of the initial signal frequency and of the detector configuration (triangular vs. two-L) on the source’s parameters estimation are also reported. An improvement of approximately one order of magnitude in the precision of the chirp mass and mass ratio can be achieved by lowering the initial frequency from 20 Hz to 2 Hz. The two-L configuration shows instead significant improvements on the inference of the source declination, due to geographical separation of the two detectors.
@@ I Introduction
Binary neutron stars are the main sources for next-generation (XG) ground-based gravitational wave (GW) detectors, such as Einstein Telescope (ET) Hild et al. (2011); Hild (2012); Abac et al. (2025); Maggiore et al. (2020); Punturo et al. (2010a) and Cosmic Explorer (CEReitze et al. (2019); Evans et al. (2021). They are expected to be observed with a rate of $7\times 10^{4}$ mergers per year Belgacem et al. (2019); Kalogera et al. (2021); Iacovelli et al. (2022). ET is designed with a xylophone configuration, combining a cryogenic interferometer for the low-frequency band with a high-power interferometer targeting frequencies above ${\sim}100~{\rm Hz}$. This design will not only significantly enhance current GW detector sensitivity, but also enable the observations of postmerger (PM) signals at kiloHertz frequencies. GWs observations of the full spectrum (inspiral-merger-postmerger) of a BNS merger can convey unique information on the nuclear matter that constitutes this compact object e.g. (Bauswein et al., 2017; Breschi et al., 2019; Radice et al., 2020; Breschi et al., 2022; Wijngaarden et al., 2022), thus improving current GW constraints after GW170817 Abbott et al. (2017a, b, 2018). While the inspiral merger (IM) carries the footprint of the matter properties at the progenitor densities, i.e. $\rho\lesssim{2}\,\rho_{\rm sat}$ (where $\rho_{\rm sat}\simeq 2.7{\times}10^{14}~{\rm g}~{\rm cm}^{-3}$ is the nuclear saturation density), the kiloHertz PM transient encodes information about the high-density (${\sim}3-6\rho_{\rm sat}$) regime of the equation of state (EoSBreschi et al. (2022), enabling the investigation of unrevealed properties of the nuclear matter at high densities, unreachable with modern experiments Radice et al. (2017); Bauswein et al. (2019); Breschi et al. (2022); Most and Raithel (2021); Prakash et al. (2021).
@@
Inference of source properties and constraints on the neutron star (NS) EoS requires the availability of complete waveform templates for matched-filtering analyses. These templates are crucially designed using empirical relations computed from numerical relativity (NR) simulations and connecting morphological features of the PM signal to NS properties, e.g. Shibata and Uryu (2002); Stergioulas et al. (2011); Bauswein and Janka (2012); Bauswein et al. (2012); Hotokezaka et al. (2013); Takami et al. (2014); Bauswein and Stergioulas (2015); Bernuzzi et al. (2015); Breschi et al. (2022, 2024a). In particular, Breschi et al. (2019) proposed the first time-domain full spectrum model that was constructed by combining an effective-one-body (EOB) IM waveform with a NR-informed analytical PM model (see also Breschi et al. (2024a, 2025)). Similar approaches have been used in Wijngaarden et al. (2022); Puecher et al. (2023a), where phenomenological frequency-domain IM waveforms are stitched to PM models based on wavelets or Lorenzian templates. In all cases the phenomenological PM parameters are connected to the source parameter using NR empirical relations; see also (Bauswein et al., 2016; Bose et al., 2018; Soultanis et al., 2022; Vretinaris et al., 2026) for other examples of analytical PM models and (Easter et al., 2020; Whittaker et al., 2022) for data-driven PM models, which have not been incorporated in full spectrum models yet.
@@
These modeling efforts, combined with the exceptional sensitivity of XG detectors, provide a crucial foundation for accurate EoS inference. Notably, even a single detection of BNS with ET may suffice to place stringent constraints on the NS mass-radius relation Breschi et al. (2022). That approach enabled tight constraints of the mass-radius diagram and demonstrated that the maximum NS mass can be determined with an accuracy better than 12%. Additional works Breschi et al. (2023); Rezzolla and Takami (2016); Chatziioannou et al. (2017); Lehoucq et al. (2025); Vretinaris et al. (2026) have similarly combined information from the IM section to PM features, highlighting the complementarity of these two regimes in probing supranuclear matter. Numerous proof of principle studies focus exclusively on the PM GW emission from BNSs to investigate the high-density nuclear matter through NR simulations. These works exploit the connection between peak frequencies and EoS-dependent properties, such as total mass and compactness Bauswein et al. (2014); Most and Raithel (2021); Easter et al. (2019); Takami et al. (2015); Bauswein et al. (2016); Ecker et al. (2025); Bamber et al. (2025); Mitra et al. (2025). Furthermore, recent works explored the potential for EoS inference using advanced and XG detectors Walker et al. (2024); Gupta et al. (2022); Iacovelli et al. (2023); Bandopadhyay et al. (2024); Kiendrebeogo et al. (2023) and assess the impact of different waveform models on nuclear EoS inference Yelikar et al. (2024).
@@
In this paper, we perform an extensive investigation of the capabilities of ET to infer BNS properties. We firstly focus on the low-frequency regime, assessing the impact on source parameter estimation of the initial frequency and of detector configurations (two-L and triangular). We then concentrate in the detections of the complete GW spectrum and study the prospective accuracy of EoS inference. We develop and employ a full spectrum model which combines the fast frequency-domain EOB TEOBResumSPA Akcay et al. (2019); Gamba et al. (2021a, 2023) with the frequency-domain wavelet-based NRPMw model Breschi et al. (2024a, 2025) (paper I and paper II hereafter). We then perform Bayesian mock analyses of 20 fiducial BNSs, covering the frequency range from $f_{0}=20$ Hz up to the kiloHertz regime. These analyses serve not only to demonstrate the feasibility of combining information across the full GW spectrum, from inspiral to merger and postmerger, but also to quantify the achievable accuracy in the inference of source parameters. Beyond parameter estimation, our primary objective is to assess the improvements in constraints on the supranuclear properties governing NS matter. By incorporating both low- and high-frequency components of the signal, we explore how analyses of ET signals enhance our ability to probe the dense-matter regime, including tighter bounds on the mass-radius relation, the maximum mass supported by a stable nonrotating NS and the PC inference.
@@
All quantities are expressed in SI units, with masses in solar masses $\rm M_{\odot}$} and distances in Mpc. The symbol $\log$ denotes the natural logarithm. The waveform strain is decomposed in multipoles $(\ell,m)$ as
@@
where $D_{L}$ is the luminosity distance, $_{-2}Y_{\ell m}$ are the s=2s=-2 spin-weighted spherical harmonics and $\iota,\varphi$ are, respectively, the polar and azimuthal angles that define the orientation of the binary with respect to the observer. Each waveform mode $h_{\ell m}(t)$ is decomposed in amplitude $A_{\ell m}(t)$ and phase $\phi_{\ell m}(t)$, as
@@
The waveform depends on intrinsic and extrinsic source parameters. The intrinsic parameters of a BNS system are masses, spins and tidal polarizability parameters, e.g. $\boldsymbol{\theta}_{\rm bin}=\{M,q,\chi_{1},\chi_{2},\Lambda_{1},\Lambda_{2}\}$. The total binary mass is indicated with $M=m_{1}+m_{2}$, the mass ratio is $q=m_{1}/m_{2}\geq 1$, and the symmetric mass ratio is $\nu=m_{1}m_{2}/M^{2}$. The dimensionless spin vectors are denoted with ${\boldsymbol{\chi}}_{i}$ for $i=1,2$ and the spin component aligned with the orbital angular momentum L are labeled as $\chi_{i}={\boldsymbol{\chi}}_{i}\cdot\textbf{L}/|\textbf{L}|$. The effective spin parameter $\chi_{\rm eff}$} is a mass-weighted, aligned spin combination, i.e.,
@@
The quadrupolar tidal polarizability parameters are defined as $\Lambda_{i}=({2}/{3})\,k_{2,i}\,C_{i}^{-5}$ for $i=1,2$, where $k_{2,i}$ and $C_{i}$ are the second Love number and the compactness of the $i$th object, respectively. The reduced tidal parameter $\tilde{\Lambda}$ is
@@
The extrinsic parameters of the source ${\boldsymbol{\theta}_{\rm ext}}=\{D_{L},\iota,\alpha,\delta,\psi,t_{\rm mrg},\phi_{\rm mrg}\}$, i.e. inclination angle $\iota$, right ascension angle $\alpha$, declination angle $\delta$, polarization angle $\psi$, time of coalescence $t_{\rm mrg}$, and phase at the merger $\phi_{\rm mrg}$, allow us to identify the location and orientation of the source. Moreover, we include the recalibration parameters $\boldsymbol{\delta}_{\rm fit}$ that account for deviations from the predictions of the EoS-insensitive relations consistently with the related theoretical uncertainties. NRPMw is parametrized by two additional sets of degrees of freedom (see paper I for a detailed discussion). The PM parameters ${\boldsymbol{\theta}_{\rm PM}^{\rm free}}=\{\phi_{\rm PM},t_{c},\alpha_{\rm drift}\}$ correspond to: the PM phase shift $\phi_{\rm PM}$ that identifies the phase discontinuity after merger; the time of collapse $t_{c}$ that characterizes the collapse of the remnant into BH; and the frequency drift $\alpha_{\rm drift}$ that accounts for linear evolution of the dominant $f_{2}$ component.
@@ II.1 Waveform model
Figure 1: Attachment between TEOBResumSPA and NRPMw for a GW template of nonspinning binary with $m_{1}=1.5~{\rm M_{\odot}}$, $m_{2}=1.4~{\rm M_{\odot}}$, $\Lambda_{1}=400$, $\Lambda_{2}=600$ and $D_{L}=40~{\rm Mpc}$. Since the model is defined in the frequency domain, we obtain $h_{22}(t)$ by performing an inverse FFT of $h_{22}(f)$.
@@
We attach the EOB model for the IM to the PM NRPMw template, focusing on the dominant $(2,2)$ mode. The information required to perform the attachment of NRPMw with TEOBResumSPA is the time and the phase at merger, respectively $t_{\rm mrg}$ and $\phi_{\rm mrg}$. These data can be extracted from the EOB IM waveform. Then, resorting linear properties of the Fourier transform, the inspiral-merger-postmerger (IMPM) waveform can be computed as the sum of the premerger template $h_{\rm EOB}$ and the PM one $h_{\rm PM}$ in the frequency-domain,
@@
We also validate the model against the $(2,2)$-mode of six NR simulations that are not included in the calibration set of NRPMw, finding that the model is $89-95\%$ faithful to NR when considering the full spectrum of frequencies and the ET-D sensitivity curve Hild et al. (2011); Hild (2012). More details on the IMPM model and its validation can be found in Appendix A.
@@ II.2 Inference
Figure 2: Amplitude spectral density (ASD) of ET in configuration D Hild et al. (2011); Hild (2012) (in black) and comparison with ASD of LIGO L1 detector, relative to GW170817 event Abbott et al. (2019a) (in gray), to show the improvement in the detection of GW signals with XG detectors. The solid colored lines represent five exemplary BNS waveform spectra with TEOBResumSPA_NRPMw template for $\mathcal{M}=1.24~{\rm M_{\odot}}$, $q=1$, varying $\tilde{\Lambda}$ and $D_{L}$ as indicated in the legend.
We perform Bayesian parameter estimation (PE) experiments by injecting and analyzing mock zero noise BNS signals with XG detectors. We employ the power spectral density (PSD) curve of the ET detector, as reported in Fig. 2, in configuration D Hild et al. (2011); Hild (2012), i.e. an equilateral triangular shape with $10~{\rm km}$ long arms. We simulate both IM-only and IMPM waveforms and recover them using the same model employed for the injection, TEOBResumSPA and TEOBResumSPA_NRPMw respectively.
@@
Bayesian analysis is based on the evaluation of the likelihood function (See Appendix B.1 for additional details), which is the most computationally expensive part. To speed up the computations, we make use of the relative binning technique Dai et al. (2018); Zackay et al. (2018) in the IM part. Relative binning consists in the precomputation of frequency-binned overlap of a fiducial waveform with the data, since the ratio of neighboring waveforms in frequency domain is smooth in the inspiral part. In particular, for the IM-only PEs we employ the relative binning method for the whole frequency range of the integration. For the IMPM cases, we utilize relative binning only up to a cutoff frequency $f_{\rm cut}=2048~{\rm Hz}$ and we perform a normal Bayesian PE without speed up after the cutoff frequency. The two likelihoods are then combined by summing their respective contributions
@@
where $\log p({\text{{d}}}_{\rm IM}|\boldsymbol{\theta},H_{\rm IM}^{\rm RB})$ is the likelihood evaluated up to $f_{\rm cut}$ with relative binning, while $\log p({\text{{d}}}_{\rm PM}|\boldsymbol{\theta},H_{\rm PM})$ is evaluated in the high-frequency section, both using TEOBResumSPA_NRPMw model. In order to investigate how sensitive the choice of $f_{\rm cut}$ is, we run 3 additional IMPM analyses with different EoSs with $f_{\rm cut}=1024~{\rm Hz}$ and compare posteriors with the ones of the analyses presented in Sec. III. In all the cases the posteriors are completely compatible, we therefore conclude that the cutoff frequency used does not have any impacts in the result of the inference. For further explanation of the relative binning implementation and a discussion of the approximations used, we refer to Appendix B.2.
II.3 EoS constraints from quasiuniversal relations
Given an estimation of the intrinsic BNS parameters, we show how XG detectors can enable precise constraints on the EoS of NSs. Inference of IM delivers posteriors of the masses of the compact objects and of the tidal polarizability parameters. The information contained in the IMPM signals, instead, allows to go beyond $m_{i}$ and $\Lambda_{i}$ and probe high-density NS properties. Breschi et al. (2022) combined the Fisher matrix approach in the IM part with the Bayesian analysis of the PM in order to obtain a tighter constraint on the mass-radius relation. Here we perform a full spectrum Bayesian PE, aiming to obtain more reliable inference, particularly in the high-density regime. The resulting pipeline constrains different quantities of interest in EoS modeling and is described in what follows.
The first step is to map the posteriors for the total mass and the peak PM frequency $f_{2}$ into posteriors of maximum density $\rho_{\rm max}^{\rm TOV}$} and radius at the maximum mass $R_{\rm max}^{\rm TOV}$}, using the NR phenomenological relations Breschi et al. (2022, 2024b)